# Neural Audio Codecs for Prompt-Driven Universal Source Separation

### v1

@@ 1 Introduction
In this manuscript, we propose CodecSep, a novel framework for text-conditioned source separation that combines the interpretability of prompt-driven separation with the compute efficiency of neural audio codecs (NACs). CodecSep is the first model to bridge NACs with universal source separation (USS), enabling flexible extraction of arbitrary sound sources based on natural language queries. By conditioning a transformer-based masker on text embeddings generated via the text branch of CLAP—a multimodal encoder trained on audio-text pairs Wu et al. (2023)—CodecSep performs state-of-the-art (SOTA) separation directly in the NAC encoder latent space, introducing semantic control while maintaining very low compute cost.
The growing demand for flexible, real-time audio processing in bandwidth-constrained or on-device scenarios calls for generalizable and computationally efficient methods. Traditional separation models tackle the challenge of disentangling sound sources from complex mixtures Vincent et al. (2018) but are typically specialized for narrow domains such as speech or music. USS broadens this scope by separating arbitrary sources without reliance on predefined labels. Within this setting, text-guided source separation has emerged as a compelling approach, enabling users to specify the target source via natural language prompts—offering greater transparency, flexibility, and accessibility in real-world applications.
Recent advances have embraced this direction by integrating textual prompts into separation pipelines. Architectures based on the Conv-TasNet encoder-masker-decoder blueprint Luo and Mesgarani (2019) have achieved strong results in domain-specific tasks Chen et al. (2020b); Subakan et al. (2021); Li et al. (2023). Systems like AudioSep Liu et al. (2024) extend this framework to the text-conditioned setting by encoding text queries with models such as BERT Devlin et al. (2019) or CLAP Wu et al. (2023) and injecting semantic information through Feature-wise Linear Modulation (FiLM) layers Perez et al. (2018). These modifications allow intermediate representations to adapt dynamically to prompt semantics, steering the model towards the desired source.
However, such models, trained with signal-level objectives like scale-invariant signal-to-distortion ratio (SI-SDR) Luo and Mesgarani (2019); Le Roux et al. (2019), are not only highly sensitive to compression artifacts but also tend to be large and computationally intensive. This limits their practicality for on-device or real-time applications where resources are constrained. As a result, they are often deployed via the cloud, introducing additional latency and infrastructure costs.
Neural audio codecs (NACs) offer a promising alternative. Models like SoundStream Zeghidour et al. (2022), Encodec Défossez et al. (2022), and DAC Kumar et al. (2023) compress audio into discrete token sequences using encoder-quantizer-decoder pipelines. Leveraging Residual Vector Quantization (RVQ), they enable adaptive bitrate control and efficient representation. These codecs are trained on diverse audio data using a combination of adversarial and perceptual losses, which promotes both fidelity and robustness under compression. The resulting token representations have proven useful in tasks such as audio generation for multimodal models Wang et al. (2024); Du et al. (2024) and language-conditioned synthesis Borsos et al. (2023); Wang et al. (2023). Building on this foundation, recent efforts like CodecFormer Yip et al. (2024b) and SDCodec Bie et al. (2024) have explored incorporating NACs into source separation pipelines. While these approaches achieve lightweight, high-fidelity performance, they remain limited to separating fixed types of sources (e.g., speech separation).
CodecSep addresses this gap by combining the computational efficiency of NACs with the flexibility of language-guided separation. It follows a NAC-based encoder-decoder architecture and integrates a transformer-based masker conditioned on CLAP-derived text embeddings. These embeddings generate FiLM parameters that modulate each transformer layer, enabling the model to selectively extract sources that align with the user’s query. Operating entirely in the codec latent space, CodecSep delivers interpretable, high-performance separation under strict computational constraints.
@@ 2 Related Work
2.1 Audio Source Separation
Notable approaches like DPTNet Chen et al. (2020b), SepFormer Subakan et al. (2021), and TDANet Li et al. (2023) follow the encoder-masker-decoder framework. In this setup, the encoder generates STFT-like latent representations. At the same time, a masker network predicts source masks, which are then applied elementwise to the encoder outputs to obtain source-specific latent representations. These estimates are further processed by a decoder, which mirrors the encoder architecture, to reconstruct the source audio. TDANet Li et al. (2023), the latest iteration, enhances this framework with a top-down attention mechanism, using global and local attention modules to hierarchically model multi-scale acoustic features. CodecSep also adopts this encoder-masker-decoder structure, extending it to operate in the NAC latent space with text guidance.
Beyond the encoder-masker-decoder paradigm, a variety of alternative architectures have been explored. Models like Wave-UNet Stoller et al. (2018) and Demucs Défossez et al. (2019, 2021) utilize encoder-decoder designs that estimate source waveforms directly, omitting the masking stage altogether. Other approaches operate in the frequency domain, notably, models like MM-DenseLSTM Takahashi et al. (2018), DCCRN Hu et al. (2020), and UNet-based architectures such as Spleeter Hennequin et al. (2020) perform separation directly in the complex STFT domain, jointly modelling magnitude and phase components. These diverse strategies highlight the broad design space of source separation, spanning both time and frequency domains, with and without masking.
2.2 Universal Source Separation
Prior work on source separation has concentrated mainly on speech or music-specific tasks. More recent advances have shifted toward USS, which aims to disentangle arbitrary sources from a mixture without assuming predefined source categories. Supervised USS models often employ Permutation Invariant Training (PIT) Yu et al. (2017); Kavalerov et al. (2019), while unsupervised approaches like Mixture Invariant Training (MixIT) Wisdom et al. (2020) learn directly from mixtures without access to clean references. However, both approaches assume a fixed maximum number of sources in the mixture and produce all outputs indiscriminately. As a result, an additional source identification step is needed post-separation to determine which of the estimated signals corresponds to the target source.
Query-Guided Source Separation (QSS) extends USS by enabling source extraction through external queries, including visual input, exemplar audio, class labels, or text. Text queries are particularly effective, offering a compact and expressive way to convey high-level semantics and encode spatial or temporal cues of the query source—all without the need for expensive reference data. Recent work such as AudioSep Liu et al. (2024) explores text-guided separation by integrating embeddings from models like BERT Devlin et al. (2019) and CLAP Wu et al. (2023), the latter of which aligns audio and text in a shared embedding space. AudioSep employs FiLM conditioning into the separation pipeline at intermediate layers of the masker. CodecSep adopts a similar strategy but operates in the NAC encoder latent space instead.
2.3 Source Separation with Neural Audio Codecs
Recent work has explored integrating neural audio codecs into source separation pipelines. CodecFormer Yip et al. (2024b) and its extension CodecFormer-EL Yip et al. (2024a) both leverage DAC as the encoder-decoder backbone, performing separation directly in the codec’s latent embedding space via a transformer-based separation module. The original CodecFormer is trained using a negative SI-SDR loss to supervise the predicted outputs. At the same time, CodecFormer-EL introduces an embedding-level objective that minimizes the distance between the separator’s output and the DAC encoder’s ground truth latent embeddings for each source. In contrast, SD-Codec Bie et al. (2024) integrates source separation more tightly into the codec itself. It is explicitly designed to disentangle mixtures into speech, music, and SFX. To achieve this, it introduces three dedicated RVQ branches—each trained to encode a specific source category. During encoding, the outputs of these domain-specific quantizers are summed across time to form a mixture-level representation, which is then decoded to reconstruct both the whole mixture and individual sources. This design enables SD-Codec to embed separation behavior directly into the codec’s structure. In this work, we repurpose the CodecFormer transformer separator into a text-conditioned mask predictor, enabling prompt-guided control over the separation process.
@@ 3 Method
CodecSep is based on the 16kHz variant of CodecFormer Yip et al. (2024b), which we adapt to enable text-driven source separation. Specifically, we modify the transformer module to function as a source mask estimator and incorporate Feature-wise Linear Modulation (FiLM) Perez et al. (2018) to inject text-based conditioning into the separation process. FiLM layers are applied to the transformer’s intermediate activations, allowing the model to adjust its internal representations based on the semantics of the input query. To produce the FiLM parameters, a lightweight query network consumes embeddings produced by a pretrained CLAP encoder Wu et al. (2023) from the input prompt.
Figure 1 (Left) illustrates our text-guided USS pipeline. Unlike AudioSep’s raw STFT-based, compute-heavy architecture, CodecSep operates on compact codec representations, offering significantly greater compute efficiency for resource-constrained edge deployment. Figure 1 (Right) illustrates how text-conditioning is integrated into the intermediate encoder layers of the transformer masker.
3.1 Model Architecture
CodecSep utilizes DAC Kumar et al. (2023) for the encoder-decoder modules and a transformer-based masker network with text conditioning to disentangle query sources within the DAC encoder representation space. We will discuss each component in detail and present how we incorporated text conditioning within the transformer masker architecture.
3.1.1 Codec Backbone
DAC Kumar et al. (2023) builds upon the Encodec Défossez et al. (2022) and SoundStream Zeghidour et al. (2022) architectures and utilizes fully convolutional encoder and decoder networks. Leaky ReLU activations in the architectures are substituted with the periodic Snake Activation function Ziyin et al. (2020) expressed as $x+\sin^{2}{x}$. This enhances the modeling of periodic audio signals by incorporating a periodic inductive bias. Residual vector quantization is integrated into the encoder output, which enables audio compression at varying bit rates. In addition, factorized codes—where code assignment occurs in a lower-dimensional space than the codebook vectors—and L2-normalized codebook vectors are utilized to enhance codebook efficiency.
Let us consider compressing an audio recording $x(t)$ of duration $1$s, sampled at $F_{s}=24$kHz to target bit-rate of R=6000R=6000bps. The encoder network $Enc(\cdot)$ down-samples $x(t)$ by a factor of $M=320$ and outputs $T=\frac{F_{s}}{M}=75$ frames of latent representations, denoted as $Z=[z_{t}\in R^{d}|t\in[T]]$. Each frame $z_{t}$ is allocated $r=\frac{R}{T}=80$ bits. Residual vector quantization (RVQ) efficiently allocates the bit-rate budget across a cascade of $N_{q}=8$ vector quantizers. The first vector quantizer in the cascade quantizes the frame $z_{t}$, and the residuals (the quantization error) are computed. The residuals are then iteratively quantized by the remaining $N_{q}-1$ quantizers in the cascade. Each quantizer is assigned $r_{i}=\frac{r}{N_{q}}=10$ bits, maintaining a manageable code-book size of $2^{10}=1024$. The number of quantizers in the cascade $N_{q}$ is key in balancing computational complexity and compression for a given bit-rate budget $r$.
The RVQ network $Quant(\cdot)$ produces a code matrix $A=[a_{t}\in[1024]^{8}|t\in[T]]$ given encoder output $Z$. Each frame $z_{t}\in Z$ is quantized into a code vector $a_{t}\in[1024]^{8}$, which can be mapped to an embedding via code-book lookup as $e_{t}=\sum_{i=1}^{8}e^{i}_{t}$, where $e^{i}_{t}$ corresponds to the embedding associated with code $a^{i}_{t}\in[1024]$ from the $i^{th}$ RVQ code-book. The RVQ output embedding sequence $E=[e_{t}|t\in[T]]$ is then processed by the decoder network $Dec(.)$, which up-samples it to generate $y(t)$, a reconstruction of the original audio $x(t)$.
3.1.2 Transformer Masker
CodecSep masker module $Mask(.)$ predicts source mask given DAC Kumar et al. (2023) encoder $Enc(.)$ representation $Z$. These masks are applied element-wise to representation $Z$ to extract the corresponding source representation estimate. It follows CodecFormer Yip et al. (2024b) Transformer separator module architecture and is designed as a stack of $L=16$ transformer layers $Mask^{l\in[L]}(.)$ of embedding size $d=256$. The output of the final transformer layer, $H^{L}=Mask^{L}(H^{L-1})$, is processed through $K$ distinct 1D convolutional layers to generate source masks $\{M_{k}|k\in[K]\}$. To preserve the periodic inductive bias across the entire model, the Snake activation function, employed in DAC Kumar et al. (2023), replaces tanh and LeakyReLU activations throughout the masker module. The resulting source representation estimates are subsequently mapped to time domain $\{\tilde{y}_{k}=Dec(M_{k}\odot Z)|k\in[K]\}$ using DAC Kumar et al. (2023) decoder $Dec(.)$ directly, without being processed by RVQ network $Quant(.)$.
3.1.3 Text-guided Source Separation
We consider a single-channel audio mixture $x(t)\in\mathbb{R}^{1\times N}$, which consists of an unknown number of overlapping open-domain sound sources. The mixture is expressed as the following sum:
$x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$ (1)
where $\mathcal{S}$ is the (potentially unbounded) set of sound source classes or instances present in the mixture. Our objective is to recover a target source waveform or stem $y_{s}(t)$ from the mixture $x(t)$, using a natural language query $\tau_{s}$ that semantically describes the desired source. These prompts may reference specific sources (e.g., “dog barking,” “violin solo,” “crowd cheering”) or combinations of sources (e.g., “speech and music,” “footsteps and glass breaking”), allowing for fine-grained control.
To integrate text-based conditioning into CodecSep, we leverage CLAP Wu et al. (2023) and utilize a masker module $Mask(.)$ with $K=1$ final convolution to generate a single source mask $M_{s}$. Given a query text $\tau$ that describes the target source, CLAP Wu et al. (2023) generates a corresponding text embedding $e_{\tau}\in\mathbb{R}^{d}$. This embedding is used to modulate the intermediate $l\in[2,...,L-1]$ transformer layers $Mask^{l}(.)$ via feature-wise linear modulation (FiLM) Perez et al. (2018), enabling the network to estimate the desired source conditioned on the text input.
At transformer layer $l$, the layer output $H^{l}\in\mathbb{R}^{d\times T}$ is modulated along the feature dimension using parameters $\gamma^{l},\beta^{l}\in\mathbb{R}^{d}$. These modulation parameters are generated by a query network $query(.)$, which processes text embedding $e_{\tau}$ to produce $(\gamma^{l},\beta^{l})_{l=1}^{L}=query(e_{\tau})$. The modulation transformation is then carried out as follows:
$\tilde{H}^{l}_{s}=FiLM(H^{l}_{i},\gamma^{l}_{i},\beta^{l}_{i})=\gamma^{l}_{i}H^{l}_{i}+\beta^{l}_{i},\forall i\in[d]$ (2)
The decoder network $Dec(.)$ subsequently processes the source representation estimate $\tilde{Z}_{s}=M_{s}\odot Z$ and generates an estimate of the query source audio $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$.
3.2 Training Objective
We train CodecSep to reconstruct each individual source from mixtures using textual prompts corresponding to broad categories such as music, speech, and a wide range of sound effects (SFX). We also reconstruct the input mixture $x$ by decoding the summed source latent estimates: $\tilde{x}=g(\sum_{s}\tilde{Z}_{s})$. The model is optimized using the negative scale-invariant signal-to-distortion ratio (SI-SDR) loss Luo and Mesgarani (2019); Le Roux et al. (2019), encouraging the faithful separation of sources aligned with the given queries.
To support generalization across varied SFX types, prompts associated with SFX are not limited to atomic events—they may describe composite scenarios involving multiple overlapping sources (e.g., “footsteps, background traffic,” “crowd cheering, fireworks”), requiring the model to isolate multi-faceted acoustic scenes based on complex queries.
SI-SDR is designed to measure the similarity between the predicted $\tilde{v}$ and ground truth $v$ signals, focusing on the signal-to-noise ratio (SNR) while being invariant to scale differences. It is given as,
@@ 3.2 Training Objective
The overall loss function is then given as:
[…]
To account for compression artifacts and improve training stability, we further compute the negative codec SI-SDR (cSI-SDR) Yip et al. (2024b), which is a variant of SI-SDR calculated with a codec-processed version of ground truth audio. During training, we keep the DAC Kumar et al. (2023) backbone and CLAP Devlin et al. (2019) text encoder weights frozen, updating only the parameters of $Mask(.)$ and $query(.)$ modules.
@@ 3.3 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
Problem setup and pipeline contrast.
Let $x(t)\in\mathbb{R}$ be a mono mixture with sources $\{y_{s}(t)|s\in\mathcal{S}\}$, where $x(t)$ is expressed as equation 1.
@@
$x\;\xrightarrow{\text{STFT}}\;X\!\in\!\mathbb{C}^{F\times T}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\odot|X|\exp(\angle X+\angle\hat{M}_{s})\;\xrightarrow{\text{ISTFT}}\;\tilde{y}_{s}(t),$
where $g(\cdot,e_{\tau})$ denotes a FiLM-conditioned, complex-domain spectrogram separator that predicts a magnitude mask $|\hat{M}_{s}|\in[0,1]^{F\times T}$ and a phase residual $\angle\hat{M}_{s}$, conditioned jointly on the mixture spectrogram $X$ (obtained by the STFT of $x$) and the text embedding $e_{\tau}$. The predicted magnitude mask and phase residual are then applied element-wise to form the source spectrogram $\tilde{Y}_{s}$ which is subsequently transformed back to the time domain via ISTFT to obtain $\tilde{y}_{s}(t)$.

### v2

@@ 1 Introduction
We propose CodecSep, a text-conditioned universal source separation (USS) framework that marries the interpretability of prompt-driven extraction with the efficiency of neural audio codecs (NACs). To our knowledge, CodecSep is the first system to bridge NACs with USS: it conditions a transformer masker on CLAP text embeddings Wu et al. (2023) via Feature-wise Linear Modulation (FiLM) Perez et al. (2018), and performs separation directly in the codec encoder latent space. This design introduces semantic control while preserving the low compute footprint of codec representations, enabling on-device and real-time deployment.
Flexible, real-time separation on bandwidth- or compute-constrained platforms remains challenging. Classic models disentangle sources from complex mixtures Vincent et al. (2018) but are often domain-specific (e.g., speech/music) and heavy. Recent text-guided systems like AudioSep Liu et al. (2024) extend encoder–masker–decoder designs (e.g., Conv-TasNet-style Luo and Mesgarani (2019)) by injecting semantics from BERT/CLAP through FiLM layers Devlin et al. (2019); Wu et al. (2023); Perez et al. (2018). However, spectrogram/waveform-domain separators trained with SI-SDR-style losses Luo and Mesgarani (2019); Le Roux et al. (2019) are compute-intensive and sensitive to compression artifacts, often pushing inference to the cloud.
NACs such as SoundStream, Encodec, and DAC Zeghidour et al. (2022); Défossez et al. (2022); Kumar et al. (2023) compress audio to discrete tokens with Residual Vector Quantization (RVQ), providing compact, perceptually aligned latents useful for generation and conditioned synthesis Borsos et al. (2023); Wang et al. (2023, 2024); Du et al. (2024). Prior codec–separation hybrids (CodecFormer, SDCodec) Yip et al. (2024b); Bie et al. (2024) demonstrate light-weight, high-fidelity separation but target fixed stems (e.g., speech separation or speech vs. music vs. SFX) rather than open-domain prompt-conditioned USS.
CodecSep adopts a frozen DAC encoder–decoder backbone and inserts a FiLM-conditioned transformer masker that predicts a soft mask over codec latents. CLAP-derived text embeddings Wu et al. (2023) are mapped to per-layer FiLM parameters, modulating the masker’s intermediate activations to align the selected latent subspace with the query semantics. Operating on compact codec features cuts memory traffic and MACs compared to spectrogram-domain pipelines, while preserving the codec’s inductive biases (periodicity, timbre, transients). In doing so, CodecSep delivers interpretable, prompt-guided separation with markedly lower compute, without sacrificing separation fidelity.
@@ 2 Related Work
Classical audio source separation systems frequently adopt an encoder–masker–decoder design in which an encoder produces STFT-like latents, a masker predicts source-specific masks, and a decoder reconstructs waveforms. Representative models include DPTNet Chen et al. (2020b), SepFormer Subakan et al. (2021), and TDANet Li et al. (2023), the last introducing a top-down attention scheme that blends global and local attention to capture multi-scale acoustic structure. Beyond masking pipelines, several works generate waveforms directly in the time domain (Wave-UNet Stoller et al. (2018), Demucs Défossez et al. (2019, 2021)) or operate fully in the complex STFT domain with joint magnitude–phase modeling (MM-DenseLSTM Takahashi et al. (2018), DCCRN Hu et al. (2020), Spleeter Hennequin et al. (2020)), underscoring the breadth of design choices across time and frequency domains.
Moving from domain-specific separation to universal source separation (USS), supervised systems often rely on Permutation Invariant Training (PIT) Yu et al. (2017); Kavalerov et al. (2019), while unsupervised approaches such as MixIT Wisdom et al. (2020) learn from mixtures alone. However, both typically assume a fixed maximum number of sources and emit all estimates indiscriminately, necessitating a separate identification step. Query-Guided Source Separation (QSS) addresses this limitation by conditioning extraction on external queries—visual, exemplar audio, class labels, or text. Text queries are compact and expressive, capture high-level semantics, and avoid costly references. AudioSep Liu et al. (2024) exemplifies this direction by injecting BERT Devlin et al. (2019) or CLAP Wu et al. (2023) embeddings via FiLM Perez et al. (2018) at intermediate masker layers to steer separation toward the requested source.
Neural audio codecs (NACs) have recently been integrated into separation pipelines to improve efficiency. CodecFormer Yip et al. (2024b) performs separation directly in DAC Kumar et al. (2023) latent space with a transformer separator trained using negative SI-SDR, and CodecFormer-EL Yip et al. (2024a) further adds an embedding-level objective that aligns separator outputs with DAC encoder latents. SD-Codec Bie et al. (2024) embeds separation behavior inside the codec by assigning dedicated RVQ branches to speech, music, and SFX, summing their codes to form mixture-level representations while enabling per-stem decoding. In contrast, our work repurposes the CodecFormer transformer into a text-conditioned mask predictor operating in the NAC latent domain: CLAP-driven FiLM conditioning modulates intermediate transformer activations to estimate a single, prompt-guided mask, enabling efficient, controllable USS without fixed source heads.
@@ 3 Method
CodecSep adapts the 16 kHz DAC Kumar et al. (2023) codec backbone for text-driven universal source separation (USS). We use a transformer masker that estimates soft masks over codec latents and inject text conditioning via Feature-wise Linear Modulation (FiLM) Perez et al. (2018) using CLAP text embeddings Wu et al. (2023). FiLM is applied to intermediate transformer activations so the query semantics steer separation. Figure 1 (Left) shows the overall text-guided pipeline; Figure 1 (Right) highlights the FiLM-conditioned masker. Compared to STFT-domain AudioSep, operating in compact codec latents yields markedly lower compute and is amenable to edge deployment.
3.1 Our Model
Backbone codec.
We use DAC Kumar et al. (2023) as encoder–decoder. Following Encodec/SoundStream, DAC uses fully-convolutional encoder/decoder with the periodic Snake activation $x+\sin^{2}x$ (replacing LeakyReLU) to bias periodic audio modeling. Residual vector quantization (RVQ) compresses encoder outputs with factorized codes and $\ell_{2}$-normalized codebooks.
For a 1 s audio $x(t)$ at $F_{s}{=}24$ kHz compressed to R=6000R{=}6000 bps, the encoder $Enc(\cdot)$ downsamples by M=320M{=}320 to $T{=}{F_{s}}/{M}{=}75$ frames $Z{=}[z_{t}\!\in\!\mathbb{R}^{d}]_{t=1}^{T}$ with r=R/T=80r{=}R/T{=}80 bits/frame. RVQ allocates $r_{i}{=}r/N_{q}{=}10$ bits across $N_{q}{=}8$ codebooks (size $2^{10}{=}1024$). Given $Z$, $Quant(\cdot)$ yields discrete codes $A{=}[a_{t}\!\in\![1024]^{8}]$, which map to embeddings $e_{t}{=}\sum_{i=1}^{8}e^{i}_{t}$; $Dec(\cdot)$ upsamples $E{=}[e_{t}]$ back to waveform $y(t)$.
Transformer masker and text-guided separation.
We consider a mono mixture as $x(t)=\sum_{s\in\mathcal{S}}y_{s}(t)$, where $x(t)$ is the observed waveform, $\mathcal{S}$ is the (unbounded) set of source classes/instances present, and $y_{s}(t)$ is the waveform of source $s$. Passing $x(t)$ through the frozen DAC encoder yields codec latents $Z\in\mathbb{R}^{d\times T}$ ($d$: channel width, $T$: latent frames). The masker $Mask(\cdot)$ operates in this latent space to predict an element-wise mask that selects the target source. We adopt a CodecFormer-style transformer with L=16L{=}16 layers, width d=256d{=}256, and Snake activations. Given a natural-language query $\tau$ (e.g., “dog barking”, “speech and music”), we compute a CLAP text embedding $e_{\tau}\in\mathbb{R}^{d}$. A lightweight query network $query(.)$ maps $e_{\tau}$ to per-layer FiLM parameters $(\gamma^{l},\beta^{l})\in\mathbb{R}^{d}$ for $l\in\{2,\ldots,L{-}1\}$, applied channel-wise to intermediate activations $H^{l}\in\mathbb{R}^{d\times T}$: $\tilde{H}^{l}=\mathrm{FiLM}(H^{l};\gamma^{l},\beta^{l})=\gamma^{l}\odot H^{l}+\beta^{l}$. The final transformer output $H^{L}$ is then passed through a single 1D convolutional head to produce the prompt-conditioned mask $M_{s}\in[0,1]^{d\times T}$. We obtain source latents $\tilde{Z}_{s}=M_{s}\odot Z$ and decode with the frozen codec decoder to get the waveform estimate $\tilde{y}_{s}=Dec(\tilde{Z}_{s})$, bypassing RVQ lookup.
Training Objective
@@ 3.2 Design Rationale: FiLM-Conditioned Masking in NAC Latent Space
$x\xrightarrow{\mathrm{STFT}}X\!\in\!\mathbb{C}^{F\times T_{spec}}\;\xrightarrow{\,g(X,e_{\tau})\,}\;\tilde{Y}_{s}=|\hat{M}_{s}|\!\odot\!|X|\,\exp\big(\angle X+\angle\hat{M}_{s}\big)\xrightarrow{\mathrm{ISTFT}}\tilde{y}_{s}(t),$
where $g(\cdot,e_{\tau})$ is a FiLM-conditioned separator that predicts a magnitude mask $|\hat{M}_{s}|\!\in\![0,1]^{F\times T_{spec}}$ and a phase residual $\angle\hat{M}_{s}$ from $X$ and text embedding $e_{\tau}$.